# Universal Prompt Optimizer for Safe Text-to-Image Generation

### v3

@@ 1 Introduction
Despite their wide adoption, T2I models are also used by malicious users to generate unsafe content like sexual, harassment and illegal-activity images Schramowski et al. (2023). Although T2I models have been developed to generate safe content through filtering training data  [^1] https://stability.ai/news/stable-diffusion-v2-release or robust learning Blau et al. (2022) in the developing stage, recent works Gao et al. (2023); Qu et al. (2023); Chin et al. (2023) have shown that T2I models are still vulnerable to prompt perturbations, which make these models disrupted to generate inappropriate content. Therefore, further defensive methods have been proposed, such as rejecting the generation of toxic images by detection Rando et al. (2022), guiding the model to generate safety content through embedding blocking Rombach et al. (2022) or fine-tuning Gandikota et al. (2023). Though methods to some extent alleviate the generation of harmful content, directly rejecting images can affect user usability, and optimizing the model would require obtaining the internal structure of T2I models and could be time-consuming, which lacks universality in real-world scenarios.
@@ 2 Related Work
However, the great ability of text-conditioned image generation ability also brings the risk of generating inappropriate/unsafe images, such as images containing pornographic or violent content. These inappropriate images may have a negative impact on society, thereby affecting people’s trust in AI technology. Therefore, some initial efforts have been taken to prevent the generation of inappropriate images from DMs. Generally, they could roughly be classified into two categories: detection-based approaches and removal-based approaches. Detection-based approaches Rando et al. (2022) detect generated images by using a safety checker and will refuse to output the image if the image is detected as problematic. Removal-based approaches can be further divided into two categories: guidance-based methods and fine-tuning-based methods. Guidance-based methods prevent the generation of certain concepts by blocking the text embedding of certain words or concepts during the inference stage, such as SD with Negative Prompts (SD-NP) Rombach et al. (2022) and Safe Late Diffusion (SLD) Schramowski et al. (2023). Fine-tuning-based methods like Erased Stable Diffusion (ESD) Gandikota et al. (2023), suppress the generation of certain concepts by fine-tuning the DM. These methods either return a black image when detecting inappropriate content, potentially upsetting users, or they need knowledge of T2I’s internal structure, lacking practical applicability. Our work is inherently different from existing works: (i) Our proposed framework prevents the generation of inappropriate images by directly and automatically optimizing prompts; and (ii) It can be applied to various T2I models without requiring knowledge of its internal structure.
@@ 4 Experiments
Figure 4: Illustration and comparison of different methods for removing inappropriate content on SD v1.4. Some images were blurred manually after generation for display purposes. For prompts see Appendix A.3.
@@ 4.1 Datasets
Firstly, we extract 50 prompts from each of the 6 categories in I2P Schramowski et al. (2023), namely sexual, harassment, self-harm, illegal activity, shocking, and violence, forming an evaluation dataset. The remaining prompts in I2P are split into two parts for the SFT and PPO stages, respectively. We employ the method in section 3.1 to process toxic prompts in I2P for SFT, to create a toxic-clean prompt pairs dataset for SFT. We also use Template prompts Qu et al. (2023) as another evaluation dataset. Template prompts is a manually created prompt dataset, where phrases are filled in a fixed prompt template. Prompts in it have a high probability of causing SD to generate inappropriate images. Table 1 summarizes these datasets.
@@ 4.2 Baseline Methods
To test the results on SD, we choose several state-of-the-art concept removal methods. For guidance-based methods, we choose SLD Schramowski et al. (2023) under 4 settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” Schramowski et al. (2023) for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, a learning rate of 1.9e-5, and a gradient accumulation step of 4. We set the $\beta$ in Eq. 4 to 0.02. We fine-tune the model for 1 epoch (52 steps in total). Similar to Prompist Hao et al. (2022), we generate 10 images per prompt and compute the average reward to reduce variance.
@@ 4.4 Evaluation Metric
We generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is classified as inappropriate if one or both of the classifiers output the respective label as in SLD Schramowski et al. (2023). Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS): We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model having an image encoder and a text encoder, and BLIP similarity is calculated based on image embedding and text embedding. We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.

### v4

@@ 1 Introduction
Despite their wide adoption, T2I models are also used by malicious users to generate unsafe content like sexual, harassment and illegal-activity images Schramowski et al. (2023). Although T2I models have been developed to generate safe content through filtering training data  [^1] https://stability.ai/news/stable-diffusion-v2-release or robust learning Blau et al. (2022) in the developing stage, recent works Gao et al. (2023); Qu et al. (2023); Chin et al. (2023); Tsai et al. (2023) have shown that T2I models are still vulnerable to prompt perturbations, which make these models disrupted to generate inappropriate content. Therefore, further defensive methods have been proposed, such as rejecting the generation of toxic images by detection Rando et al. (2022), guiding the model to generate safety content through embedding blocking Rombach et al. (2022) or fine-tuning Gandikota et al. (2023). Though methods to some extent alleviate the generation of harmful content, directly rejecting images can affect user usability, and optimizing the model would require obtaining the internal structure of T2I models and could be time-consuming, which lacks universality in real-world scenarios.
@@ 2 Related Work
However, the great ability of text-conditioned image generation ability also brings the risk of generating inappropriate/unsafe images, such as images containing pornographic or violent content. These inappropriate images may have a negative impact on society, thereby affecting people’s trust in AI technology. Therefore, some initial efforts have been taken to prevent the generation of inappropriate images from DMs. Generally, they could roughly be classified into two categories Tsai et al. (2023): detection-based approaches and removal-based approaches. Detection-based approaches Rando et al. (2022) detect generated images by using a safety checker and will refuse to output the image if the image is detected as problematic. Removal-based approaches can be further divided into two categories Chin et al. (2023): guidance-based methods and fine-tuning-based methods. Guidance-based methods prevent the generation of certain concepts by blocking the text embedding of certain words or concepts during the inference stage, such as SD with Negative Prompts (SD-NP) Rombach et al. (2022) and Safe Late Diffusion (SLD) Schramowski et al. (2023). Fine-tuning-based methods like Erased Stable Diffusion (ESD) Gandikota et al. (2023), suppress the generation of certain concepts by fine-tuning the DM. These methods either return a black image when detecting inappropriate content, potentially upsetting users, or they need knowledge of T2I’s internal structure, lacking practical applicability. Our work is inherently different from existing works: (i) Our proposed framework prevents the generation of inappropriate images by directly and automatically optimizing prompts; and (ii) It can be applied to various T2I models without requiring knowledge of its internal structure.
@@ 4 Experiments
Figure 4: Illustration and comparison of different methods for removing inappropriate content on SD v1.4. Some images were blurred manually after generation for display purposes. The prompts and their corresponding seeds and guidance scales are in Appendix A.3.
@@ 4.1 Datasets
Firstly, we extract 50 prompts from each of the 6 categories in I2P Schramowski et al. (2023), namely sexual, harassment, self-harm, illegal activity, shocking, and violence, forming an evaluation dataset. The remaining prompts in I2P are split into two parts for the SFT and PPO stages, respectively. We employ the method in section 3.1 to process toxic prompts in I2P for SFT, to create a toxic-clean prompt pairs dataset for SFT. We also use Template prompts Qu et al. (2023) as another evaluation dataset. Template prompts is a manually created prompt dataset, where phrases are filled in a fixed prompt template. Prompts in it have a high probability of causing SD to generate inappropriate images. Table 1 summarizes these datasets. Note that we originally select 3,561 prompts from I2P for SFT, but due to some prompts being rejected by GPT, we end up with a total of 3,318 pairs.
@@ 4.2 Baseline Methods
To test the results on SD, we choose several concept removal methods that can remove inappropriate content. For guidance-based methods, we choose SLD Schramowski et al. (2023) under 4 settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with the strength of negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” Schramowski et al. (2023) for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, a learning rate of 1.9e-5, and a gradient accumulation step of 4. We set the $\beta$ in Eq. 4 to 0.02. We fine-tune the model for 1 epoch (52 steps in total). We choose the model with the highest reward in the PPO phase. Similar to Promptist Hao et al. (2024), we generate 10 images per prompt and compute the average reward to reduce variance. Our code is built based on LLaMA-Factory Zheng et al. (2024).
@@ 4.4 Evaluation Metric
We generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is classified as inappropriate if one or both of the classifiers output the respective label as in SLD Schramowski et al. (2023). Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS). We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model. The BLIP similarity is calculated based on image embedding and text embedding Qu et al. (2023). We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.